# Log-majorizations between quasi-geometric type means for matrices

### v1

@@ 2 Preliminaries
Several basic facts on the quasi matrix means defined in (i)–(vii) have been collected, though restricted to the case $0<\alpha<1$, in Section 2.1 of our previous paper [33]. A few of those are supplemented by the next proposition and theorem including the case $\alpha>1$.
@@ 6 Concluding remarks
(2) Another topic we have addressed is the joint concavity/convexity of the trace functions $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for each $\mathcal{M}$ from $R,G,SG,\widetilde{S}G,LE$. The cases of $\mathrm{Tr}\,R_{\alpha,p}$ ($\alpha>0$) and $\mathrm{Tr}\,G_{\alpha,p}$ ($0<\alpha<1$) are already known [53, 30]. In Section 5 we have aimed at characterizing the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for unknown cases. Our method is based on the close connection between the joint concavity/convexity of geometric type trace functions and the monotonicity under CPTP maps of the generated Rényi type divergences described in Theorem 5.3. Appealing to theory of quantum $\alpha$-Rényi divergences we have obtained a necessary condition for $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ to be jointly concave/convex. The condition is the sandwiched inequalities of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ with $\mathrm{Tr}\,R_{\alpha,1/\alpha}$ and $\mathrm{Tr}\,G_{\alpha,1}$ (see Theorem 5.8), which are related to the log-majorizations discussed in Section 3. However, we cannot obtain a sufficient condition for the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ in this way, while the necessary condition obtained can be best possible, for example, as exemplified in Remark 5.10. Thus the problem is still left open for $\mathrm{Tr}\,G_{\alpha,p}$ ($\alpha>1$), $\mathrm{Tr}\,SG_{\alpha,p}$ and $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ as mentioned in Problems 5.18 and 5.20.
@@
(4) Appendix C below is a supplement to Theorems 5.3 and 5.8. It clarifies the structure of quantum $\alpha$-Rényi type divergences that satisfy the monotonicity under a subclass of quantum channels consisting of quantum-classical and classical-quantum channels. At the moment we have no explicit example of quantum $\alpha$-Rényi divergence that is not monotone under all quantum channels but monotone under the above subclass. Although Appendix C has not been used in the main body of this paper, it might be of independent interest from the point of view of quantum information.
@@ Appendix A Equality case in norm inequality for $G_{\alpha,p}$
Below we compute the coefficients up to the $4$th of the Taylor expansion in $t$ of the LHS of (A.2). From the Taylor expansions of $e^{tH}$ and $e^{-tK/2}$ one can write
@@ Appendix C Monotonicity under restricted CPTP maps
$D^{Q_{\alpha}}(B\|A)={1\over m}D^{Q_{\alpha}}(B^{\otimes m}\|A^{\otimes m})\geq{1\over m}D^{Q_{\alpha}}(\mathcal{E}_{A^{\otimes m}}(B^{\otimes m})\|A^{\otimes m})={1\over m}D_{\alpha}^{\mathrm{cl}}(\mathcal{E}_{A^{\otimes m}}(B^{\otimes m})\|A^{\otimes m})$
@@
$D^{Q_{\alpha}}(\Phi(b)\|\Phi(a))\leq D^{G_{\alpha,1}}(\Phi(b)\|\Phi(a))\geq D_{\alpha}^{\max}(\Phi(b)\|\Phi(a)).$
Moreover, we have $D_{\alpha}^{\max}(\Phi(b)\|\Phi(a))\leq D_{\alpha}^{\mathrm{cl}}(b\|a)=D^{Q_{\alpha}}(\mathrm{diag}(b)\|\mathrm{diag}(a))$ by (5.2) and (C.1), so that $D^{Q_{\alpha}}(\Phi(b)\|\Phi(a))\leq D^{Q_{\alpha}}(\mathrm{diag}(b)\|\mathrm{diag}(a))$, implying (ii).
@@
Therefore, it follows from (5.2) and Theorem 5.7(1) that $D^{Q_{\alpha}}(B\|A)\leq D_{\alpha}^{\max}(B\|A)=D^{G_{\alpha,1}}(B\|A)$, implying (i). ∎
@@
(1) As for $\mathrm{Tr}\,R_{\alpha,p}$, by Theorems 3.1(a) and 3.2 note that for $0<\alpha<1$, $\mathrm{Tr}\,G_{\alpha,1}\leq\mathrm{Tr}\,R_{\alpha,p}\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}$ if and only if $p\leq 1/\alpha$, and that for $\alpha>1$, $\mathrm{Tr}\,R_{\alpha,1/\alpha}\leq\mathrm{Tr}\,R_{\alpha,p}\leq\mathrm{Tr}\,G_{\alpha,1}$ if and only if $\max\{\alpha/2,\alpha-1\}\leq 1/p\leq\alpha$. Theorem C.1 says that if either of these holds, then $D^{R_{\alpha,p}}$ ($D_{\alpha,z}$ with $z=1/p$) is monotone under all semi-classical channels, and vice versa if $\alpha\geq 1/2$. Thus, in view of Theorem 5.9 we see that for any $\alpha\in[1/2,\infty)\setminus\{1\}$ the $\alpha$-$z$-Rényi divergence $D_{\alpha,z}$ satisfies the monotonicity under quantum channels if and only if it does under semi-classical channels.
(2) As for $\mathrm{Tr}\,G_{\alpha,p}$, by Proposition 4.2 and Theorem 3.2 note that for $0<\alpha<1$, $\mathrm{Tr}\,G_{\alpha,1}\leq\mathrm{Tr}\,G_{\alpha,p}\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}$ if and only if $p\leq 1$, and that for $1\leq\alpha\leq 2$, $\mathrm{Tr}\,R_{\alpha,1/\alpha}\leq\mathrm{Tr}\,G_{\alpha,p}\leq\mathrm{Tr}\,G_{\alpha,1}$ if and only if $\max\bigl\{1/2,{\alpha-1\over\alpha}\bigr\}\leq p\leq 1$. Theorem C.1 says that either of these holds if and only if $D^{G_{\alpha,p}}$ satisfies the monotonicity under semi-classical channels. Of course, this is only if $D^{G_{\alpha,p}}$ satisfies the monotonicity under quantum channels; see Theorem 5.15 and Proposition 5.16.
@@
Let $Q_{\alpha}$ be as above and assume that it satisfies, in addition to (d), (e), properties (b), (c) in Section 5.1. Let $(A_{i},B_{i})\in\mathrm{dom}\,Q_{\alpha}$, $i=1,2$, and $0<\lambda<1$. Assume that $\lambda A_{1}+(1-\lambda)A_{2}$ and $\lambda B_{1}+(1-\lambda)B_{2}$ commute. If $0\leq\alpha<1$ and condition (ii) (or (i)) of Theorem C.1(1) holds, then
@@
If $1<\alpha<\infty$ and condition (ii) (or (i)) of Theorem C.1(2) holds, then
@@
For instance, if $\alpha>1$ and $p\geq\max\bigl\{1/2,{\alpha-1\over\alpha}\}$, then Theorem 3.2(2) implies that $\mathrm{Tr}\,G_{\alpha,p}(A,B)\geq\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$. Hence in this case, $\mathrm{Tr}\,G_{\alpha,p}$ satisfies the ‘conditional’ joint convexity as in the above corollary, supplementing Proposition 5.16(1).

### v2

@@ 2 Preliminaries
Several basic facts on the quasi matrix means defined in (i)–(vii) have been collected, though restricted to the case $0<\alpha<1$, in Section 2.1 of our previous paper [33]. A few more facts are added in the next proposition and theorem including the case $\alpha>1$.
@@ 6 Concluding remarks
(2) Another topic we have addressed is the joint concavity/convexity of the trace functions $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for each $\mathcal{M}$ from $R,G,SG,\widetilde{S}G,LE$. The cases of $\mathrm{Tr}\,R_{\alpha,p}$ ($\alpha>0$) and $\mathrm{Tr}\,G_{\alpha,p}$ ($0<\alpha<1$) are already known [53, 30]. In Section 5 we have aimed at characterizing the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for unknown cases. Our method is based on the close connection between the joint concavity/convexity of geometric type trace functions and the monotonicity under CPTP maps of the generated Rényi type divergences described in Theorem 5.3. Appealing to theory of quantum $\alpha$-Rényi divergences we have obtained a necessary condition for $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ to be jointly concave/convex. The condition is the sandwiched inequalities of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ with $\mathrm{Tr}\,R_{\alpha,1/\alpha}$ and $\mathrm{Tr}\,G_{\alpha,1}$ (see Theorem 5.8), which are connected with the log-majorizations discussed in Section 3. However, we cannot obtain a sufficient condition for the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ in this way, while the necessary condition obtained can be best possible, for example, as exemplified in Remark 5.10. Thus the problem is still left open for $\mathrm{Tr}\,G_{\alpha,p}$ ($\alpha>1$), $\mathrm{Tr}\,SG_{\alpha,p}$ and $\mathrm{Tr}\,\widetilde{S}G_{\alpha,p}$ as mentioned in Problems 5.18 and 5.20.
@@
(4) Related to Theorems 5.3 and 5.8, in Appendix C below we clarify the structure of quantum $\alpha$-Rényi type divergences that satisfy the monotonicity under a subclass of quantum channels consisting of quantum-classical and classical-quantum channels. At the moment we have no explicit example of quantum $\alpha$-Rényi type divergence that is not monotone under all quantum channels but monotone under the above subclass. Although Appendix C has not been used in the main body of this paper, it might be of independent interest from the point of view of quantum information.
@@ Appendix A Equality case in norm inequality for $G_{\alpha,p}$
Below we compute the coefficients up to the fourth of the Taylor expansion in $t$ of the LHS of (A.2). From the Taylor expansions of $e^{tH}$ and $e^{-tK/2}$ one can write
@@ Appendix C Monotonicity under restricted CPTP maps
$D^{Q_{\alpha}}(B\|A)={1\over m}D^{Q_{\alpha}}(B^{\otimes m}\|A^{\otimes m})\geq{1\over m}D^{Q_{\alpha}}(\mathcal{E}_{A^{\otimes m}}(B^{\otimes m})\|A^{\otimes m})={1\over m}D_{\alpha}^{\mathrm{cl}}(\mathcal{E}_{A^{\otimes m}}(B^{\otimes m})\|A^{\otimes m}),$
@@
Therefore, it follows from (5.2) and Theorem 5.7(1) and (2) that $D^{Q_{\alpha}}(B\|A)\leq D_{\alpha}^{\max}(B\|A)\leq D^{G_{\alpha,1}}(B\|A)$, implying (i).
@@
$D^{Q_{\alpha}}(\Phi(b)\|\Phi(a))\leq D^{G_{\alpha,1}}(\Phi(b)\|\Phi(a))=D_{\alpha}^{\max}(\Phi(b)\|\Phi(a)).$
Moreover, we have $D_{\alpha}^{\max}(\Phi(b)\|\Phi(a))\leq D_{\alpha}^{\mathrm{cl}}(b\|a)=D^{Q_{\alpha}}(\mathrm{diag}(b)\|\mathrm{diag}(a))$ by (5.2) and (C.1), so that $D^{Q_{\alpha}}(\Phi(b)\|\Phi(a))\leq D^{Q_{\alpha}}(\mathrm{diag}(b)\|\mathrm{diag}(a))$, implying (ii). ∎
@@
(1) As for $\mathrm{Tr}\,R_{\alpha,p}$, by Theorems 3.1(a) and 3.2 note that for $0<\alpha<1$, $\mathrm{Tr}\,G_{\alpha,1}\leq\mathrm{Tr}\,R_{\alpha,p}\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}$ if and only if $p\leq 1/\alpha$, and that for $\alpha>1$, $\mathrm{Tr}\,R_{\alpha,1/\alpha}\leq\mathrm{Tr}\,R_{\alpha,p}\leq\mathrm{Tr}\,G_{\alpha,1}$ if and only if $\max\{\alpha/2,\alpha-1\}\leq 1/p\leq\alpha$. Theorem C.1 says that if either of these holds, then $D^{R_{\alpha,p}}$ ($D_{\alpha,z}$ with $z=1/p$) is monotone under all semi-classical channels, and vice versa if $\alpha\geq 1/2$. Thus, by Theorem 5.9 we see that for any $\alpha\in[1/2,\infty)\setminus\{1\}$ the $\alpha$-$z$-Rényi divergence $D_{\alpha,z}$ satisfies the monotonicity under quantum channels if and only if it does under semi-classical channels.
(2) As for $\mathrm{Tr}\,G_{\alpha,p}$, by Proposition 4.2 and Theorem 3.2 note that for $0<\alpha<1$, $\mathrm{Tr}\,G_{\alpha,1}\leq\mathrm{Tr}\,G_{\alpha,p}\leq\mathrm{Tr}\,R_{\alpha,1/\alpha}$ if and only if $p\leq 1$, and that for $1<\alpha\leq 2$, $\mathrm{Tr}\,R_{\alpha,1/\alpha}\leq\mathrm{Tr}\,G_{\alpha,p}\leq\mathrm{Tr}\,G_{\alpha,1}$ if and only if $\max\bigl\{1/2,{\alpha-1\over\alpha}\bigr\}\leq p\leq 1$. Theorem C.1 says that either of these holds if and only if $D^{G_{\alpha,p}}$ satisfies the monotonicity under semi-classical channels. Thus, by Theorem 5.15 we see that for $0<\alpha<1$ the monotonicity of $D^{G_{\alpha,p}}$ under quantum channels is equivalent to that under semi-classical channels, while this is not known for $\alpha>1$ (in view of Proposition 5.16(1)).
@@
Let $Q_{\alpha}$ be as above and assume that it satisfies, in addition to (d), (e), properties (b), (c) in Section 5.1. Let $(A_{i},B_{i})\in\mathrm{dom}\,Q_{\alpha}$, $i=1,2$, and $0<\lambda<1$. Assume that $\lambda A_{1}+(1-\lambda)A_{2}$ and $\lambda B_{1}+(1-\lambda)B_{2}$ commute. If $0<\alpha<1$ and condition (ii) (or (i)) of Theorem C.1(1) holds, then
@@
If $1<\alpha<\infty$ and condition (ii) (equivalently (i)) of Theorem C.1(2) holds, then
@@
For instance, if $\alpha>1$ and $p\geq\max\bigl\{1/2,{\alpha-1\over\alpha}\}$, then Theorem 3.2(2) implies that $\mathrm{Tr}\,G_{\alpha,p}(A,B)\geq\mathrm{Tr}\,R_{\alpha,1/\alpha}(A,B)$ for all $A,B\geq 0$ with $s(A)\geq s(B)$. Hence in this case, $\mathrm{Tr}\,G_{\alpha,p}$ satisfies the ‘conditional’ joint convexity by the above corollary, supplementing Proposition 5.16(1).